UT Feb 28 2021




Night Partition

  • Night split with LBTI - night center 00:32 - handover 00:27 to LBTI
  • Moon is up effectively all night (98.4%)
  • 1st half D/E-time
    • DONETwilight (19:15hr)-22:15 - HD 33632 - LBTI w/ALES
    • ALERT!MODS Field Correction check for IT 8315 - 22:30-00:27 - MODS (High winds moved in so could not complete)
    • If ALES observations can't be done, then MODS science can fill in (see below)
  • 2nd half LBTI
NO LUCI-1 SOUL AO

Night Plan for D/E time

  • Engineering Tasks
    • MODS Field Correction check for IT 8315
      • 1-1.5 hours in seeing better than 1.4 arcsec, clouds don't matter much.
      • priority=medium (problem may have been with us for a year)
      • run IDL task field_collect (same as measuring field aberrations: ComTools#Field_Data_Collection)
        • do LDG and RDG in parallel (two IDL windows)
        • Use 6-8 stars (some fields only have 6 stars; start the list a second time, if you are under an hour)
        • Stone_G is well placed at RA=10:17, and other Stone / BS fields are available if earlier or later.
  • Science
    • HD33632 brown dwarf companion observation - LBTI/ALES - (Stone) - 3hr
      • 05 13 17 +37 20 14
      • Directors time from December, lost to weather and M3 issue
      • Target from twilight + 3hrs
      • Weather: clear, <1.2-ish arcsec seeing, low water vapor
    • Nova - 0.5hr MODS-Bino (Wagner)
      • Any Conditions
      • Pending check by PI on data collected in poor conditions
    • mid-z ULIRGS
      • Seeing <1.2", thin clouds
      • FF0234-0139 - hits airmass ~2 at ~19:50 - even a single 30min exposure per night is OK
      • FF0937+3857 - 2 hr w/overheads - up all first half
      • script location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/Rothberg_Science/mid-z-ULIRGs
    • Local Merger min 0.5hr MODS-Bino (Rothberg)
      • Any Conditions
      • Can be used as a filler if needed
      • at least 1x1800sec exposure in all 4 MODS channel - but red is the highest science priority (in case something breaks...)
      • UGC 4079
      • NGC 2744
      • script location: Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/Rothberg_Science/Local_Mergers/scripts

Overview

20210227 23:42 Waking MODS and taking simSnap. Starting up LBTI WFS and all other toys to be used tonight.

20210228 0:18 Last night we had a ton of dumpSaved messages and process dumped skipped alerts streaming through the adsec scontrol gui. They continue. Resting did not resolve and Greg suggested restarting things. adsc stop/start, flat set with 639 modes

20210228 0:35 David has completed his test preset and telescope check out. Taking some MODS cals -
  • mods 1.0" flats dual grating.
  • 3K biases
  • 8K biases 20210228 Winds are high but within limits. David is opening
LBTI log:https://docs.google.com/spreadsheets/d/1NLjq6WSJSXHLfO8VKsPOhHgqU2T5KFzDFRHNqQglnxs/edit?usp=sharing

20210228 1:35. WFS went off line during preset to Capella. WFS hardware gui showed all offline and WFS arb showed offline. Tried to top and start WFS arb but things did not come up clean so just did a w_stop_all_w_start_all. Once up, OCAM not live. OCAM not live. Rebinning not working. CCD47 then went not live after a bit but after a rebin came online. Cycled BCU Interface, Slope Computer Ctrl, OCAM2k controller, master Diagnostic. No luck still. Cycle of BCU2k, then rebin x2 worked. But the processes on the hardware gui are flashing like its 1999. Comm is questionable.

And I ripped... I accidentally left the FLAO control in AUTO. During the closed loop we were off enough that the vignette caused too high forces. Closed manually and all was better.

20210228 2:15 Preset to HD33642. Block error (yet another comm glitch, because we need more of those in our lives) IT 7991, preventing use of optimize gain tonight.

20210228 2:42. While in an off position I was tweaking gains and the pupils went from nice, pretty and closed to appearing to be open. The AOS, WFS Control GUI and Adsec Control GUI all reflected that we were closed and happy. No safe skip frames shown in the Adsec control gui, but the safe skip from counter was incrementing rapidly in the mirror gui. Unfortunately did not make a note of the WFSFrame counter of MirrFrame Counter but did take Optical loop data.

AdSecArbitrator _34458|ERR| 1276|2021-02-28 02:41:18.134808| IDL > Command returns error : Timeout error
AdSecArbitrator_34458|INF| 1277|2021-02-28 02:41:18.332742| COMMANDHANDLER > Command execution failed: Fsm forced to Failure
AdSecArbitrator_34458|ERR| 1278|2021-02-28 02:41:18.332962| MAIN > LoadGainTag: FATAL - Timeout error
AdSecArbitrator_34458|ERR| 1289|2021-02-28 02:42:41.518888| ADAM-MODBUS > [TcpReceiveException] Error receiving data (errno=0, error message means: Success, received Bytes = 0 ), socket closed. (code -15003) TcpConnection receive error File TcpConnection.cpp line 133

20210228 3:48 3:55 I saw a flash of a pie shape on the mirror gui...Calling Greg & Guido. (3 additional flashed while on the line with them) Apparently crate 1 spews out bogus data periodically. This is not real or an issue. Will add to IT soon IT7566

20210228 4:58 Slewing to calibrator. Optimize gain is working again. Seeing 0.9". High winds ~20"

20210228 05:10 Running a bit behind. Still need to wavelength calibrate. Should be done in about 20 min.

20210228 5:40 Calibrator complete and reconfiguring to MODS...but the winds are over the limit so we are closing.

20210228 7:17 We had so many issues during the first half. We are going to restart the adsc and wfs softawre and start from scratch. Loading program on adsec was exceptionally slow!

Error during command LoadProgram ():

7:38 WARNING -adscecarb: illegal command for state ready (load program).
verified disk is not full. Just doing another adsc stop start. Did notice a few messages with :TCS Receive Exception Error... could that be related? Second time is a winner

-- BarryRothberg - 26 Feb 2021
Topic revision: r6 - 28 Feb 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback